### v1

# Uncertainty Distillation: Teaching Language Models to Express Semantic Confidence

Abstract
As large language models (LLMs) are increasingly used for factual question-answering, it becomes more important for LLMs to have the capability to communicate the likelihood that their answer is correct. For these verbalized expressions of uncertainty to be meaningful, they should reflect the error rates at the expressed level of confidence. However, when prompted to express confidence, the error rates of current LLMs are inconsistent with their communicated confidences, highlighting the need for uncertainty quantification methods. Many prior methods calculate lexical uncertainty, estimating a model’s confidence in the specific string it generated. In some cases, however, it may be more useful to estimate semantic uncertainty, or the model’s confidence in the answer regardless of how it is verbalized. We propose a simple procedure, uncertainty distillation, to teach an LLM to verbalize calibrated semantic confidences. Using held-out data to map initial uncertainty estimates to meaningful probabilities, we create examples annotated with verbalized probabilities for supervised fine-tuning. We demonstrate our method yields verbalized confidences that correlate with observed error rates with a small fine-tuned language model as well as with larger instruction-tuned models, and find that our semantic uncertainty correlates well with lexical uncertainty on short answers.
Uncertainty Distillation: Teaching Language Models to Express Semantic Confidence
Sophia Hager, David Mueller, Kevin Duh, Nicholas Andrews Johns Hopkins University shager2@jh.edu {dam,noa}@jhu.edu kevinduh@cs.jhu.edu
1 Introduction
Advances in LLM research have led to instruction-tuned generative models with impressive capabilities on many challenging tasks (OpenAI et al., 2024; Jiang et al., 2023; Dubey et al., 2024). While the flexibility and quality of these models is appealing, they may still hallucinate or give incorrect answers (Rawte et al., 2023; Bai et al., 2024). However, language models do not readily provide an interpretable measure of a model’s likelihood of correctness. LLMs tend to produce poorly-calibrated confidences when prompted to do so, with a propensity for being confidently incorrect Xiong et al. (2024). Furthermore, the elicited confidences may be impacted in unexpected ways by the choice of prompt (Sclar et al., 2023), such as the interpretation of “very confident” being dependent on the wording of the prompt.
There are several other approaches that can be used in lieu of prompting. Models’ token-level probabilities can be used to provide information as a measure of lexical uncertainty, which gives information about the likelihood of a generated string. This is often useful; however, the same fact can be expressed in any number of ways—“Berlin’s the capital of Germany” or “The capital of Germany is Berlin!” or “Die Hauptstadt Deutschlands ist Berlin”—all capturing the same meaning Kuhn et al. (2023). Semantic uncertainty is therefore challenging to capture, as token-level probabilities are influenced by the phrasing of an answer just as much as the semantics of the answer itself. This issue is exacerbated in models employing large vocabularies such as multilingual language models, in language models employing byte or character-level tokenization, or when using LLMs that are prone to producing extraneous outputs Xue et al. (2021); Wang et al. (2024).
Figure 1: An overview of how our method might be applied to a simple QA example. In Monte Carlo sampling with semantic normalization, we sample repeatedly from our language model, and use a normalization function to consolidate answers with the same semantic meaning. By consolidating the counts, we obtain a Monte Carlo estimate of each answer’s probability. In post-hoc calibration, we pass this estimate through a learned post-hoc calibration function to better align it with its likelihood of correctness. Finally, in self-annotation and fine-tuning, we translate these probabilities to verbalized signifiers and fine-tune a model to output verbalized confidences in addition to the answer.
We present uncertainty distillation, a scheme for fine-tuning a language model to verbalize uncertainty based on its own internal state. Notably, this is an estimate of semantic rather than lexical uncertainty, as the confidence is estimated from its internal representation without using a separate uncertainty quantification model or information from token-level probabilities. At inference time, models trained using uncertainty distillation efficiently generate a well-calibrated and interpretable statement of confidence in their answers, such as “Berlin is the capital of Germany [high confidence].” [^1] The uncertainty could be expressed in a variety of ways, including using special characters or numeric values. In principle, this approach has several benefits: for instance, semantically equivalent but lexically different predictions can be assigned the same confidence, and a single generation with multiple associated claims can each be assigned different confidences.
Rather than specializing to a specific model or task, uncertainty distillation automatically self-annotates any desired QA dataset with the chosen model’s confidence scores, which may then be used to fine-tune that model to produce verbalized confidences. At a high level, our approach consists of three steps: (1) obtaining semantic uncertainty estimates from the model on held-out calibration data; (2) post-hoc calibrating these into meaningful probabilities; and (3) teaching the model via continued training to output verbalized confidences along with its predictions.
Summary of contributions
We propose uncertainty distillation, a simple and effective scheme which uses supervised fine-tuning to teach LLMs to output calibrated semantic confidence statements along with predictions. We demonstrate that uncertainty distillation achieves well-calibrated results, both with a single-task fine-tuned model and with instruction-tuned models without initial task-specific training. We compare uncertainty distillation to a powerful lexical uncertainty quantification baseline on simple tasks, and find that we achieve similarly positive results despite operating without lexical uncertainty data at inference time. Finally, we analyze the impact of choices made at different steps of uncertainty distillation, and discuss our recommendations for using uncertainty distillation in practical applications.
2 Related Work
Linguistic calibration and verbalized confidences
Generally, calibration refers to the concept that predicted probabilities should align with the probability of correctness (Guo et al., 2017). Mielke et al. (2022) additionally propose the conception of “linguistic calibration”—that models demonstrate uncertainty or doubt through natural language when they are incorrect, determining this uncertainty by using a predictor to determine the likelihood that an answer is correct and considering that to be the model’s uncertainty. There are significant advantages to verbalizing uncertainty: for one, there is relatively low computational overhead to generate several extra tokens, while using a separate calibration model to estimate confidence and then communicate this information to the user requires more computation at inference time Yang et al. (2024). Verbalized confidences are also easily interpretable to an average user regardless of experience or background.
Lexical uncertainty quantification
Lexical uncertainty quantification metrics using information from token-level probabilities are commonly used and frequently effective Hu et al. (2023); Malinin and Gales (2021). These probabilities are easily obtainable, do not require additional inference-time compute to generate, and often provide sufficient information for downstream use cases: e.g. error correction in chain of thought Yin et al. (2024), hallucination detection Arteaga et al. (2024), or out-of-distribution data detection Hendrycks et al. (2020). However, there are several disadvantages to lexical uncertainty quantification: it relies on model probabilities which may not be well-calibrated Guo et al. (2017), and is often ineffective on calculating uncertainty of long generations Zhang et al. (2024). The latter in particular may present problems for end users, as models trained using Reinforcement Learning from Human Feedback (RLHF) may be incentivized to produce long outputs Singhal et al. (2024). It is therefore important to consider uncertainty quantification methods that do not rely on token-level probabilities to estimate uncertainty.
Semantic uncertainty quantification
In contexts where lexical uncertainty falls short, a natural method to obtain verbalized confidences might be to simply prompt a model to output confidences. This would theoretically provide an estimate of uncertainty without explicitly using token-level probabilities to estimate confidence. However, in practice, LLMs tend to overestimate their own confidence, possibly because human annotators tend to prefer texts with fewer markers of uncertainty Zhou et al. (2024). This, in turn, suggests while simply altering prompts may result in improved confidence estimates (Xiong et al., 2024; Tian et al., 2023), models may be fundamentally limited in their ability to acknowledge uncertainty without further training.
Running multiple steps at inference time may provide a better estimate of semantic probability. Xiong et al. (2024) investigate several inference-time strategies which use multiple steps to estimate model uncertainty, such as sampling several answers on the same question or noting if a model changes its answer when prompted with a misleading alternative. While these methods do lead to improvements in LLM calibration, no single intervention consistently emerges as the most successful, and the authors note there is significant scope for improvement. Kuhn et al. (2023) and Farquhar et al. (2024) more explicitly relate this to semantic uncertainty, and find that sampling $m$ predictions from the model and clustering by semantic equivalence results in a robust measure of semantic uncertainty that compares favorably to lexical uncertainty. A major disadvantage of these sampling-based approaches is their increased computational complexity at inference time, however; for instance, Kuhn et al. (2023) recommend sampling at least 10 outputs per question [^2] We compare this approach to our own in Appendix A and find that the additional computation does not consistently outperform uncertainty distillation..
3 Method
We propose a simple training recipe, illustrated in Figure 1 and described below, to allow a language model to express confidences that correlate with expected error rates on held-out data.
3.1 Monte Carlo sampling with semantic normalization
Assuming input $x$ and output $y$, we are looking to find $P(y\in Y_{equivalent}\mid x)$, the model’s likelihood of producing this answer or one that is semantically equivalent; however, this would require marginalization over an infinite set of strings $Y$. To make this a tractable problem, we use a Monte Carlo approximation, where our estimate of the models’ predictive distribution improves with $N$, at the expense of additional computational cost. Note however that we do not assume this quantity is a meaningful probability out-of-the-box due to potential overfitting or underfitting of the base model. To diagnose potential miscalibration of the base model as well as correct for it, we may use held-out calibration data that was not seen during training. [^3] We examine the efficacy of our method when this assumption does not hold true in §missing 6.1.
In more detail, to fit a post-hoc calibrator, we need a supervised dataset of datapoints not seen at training time $\{X^{\text{cal}},Y^{\text{cal}}\}$. For each example $x\in X^{\text{cal}}$ we sample $N$ candidate answers $\{\hat{y}_{i}\}_{i=1}^{N}\sim P_{\theta}(Y\mid X=x)$ from a model’s predictive distribution [^4] This model may have been fine-tuned on the specific task as in §missing 4.1 or instruction-tuned as in §missing 4.2.. Before calculating the relative frequency of strings, we apply a normalization function (or set of normalization functions) to consolidate semantically similar outputs. In the short-form QA tasks we consider in §missing 4, we use the simple normalization function of removing punctuation and standardizing capitalization, but for more complex questions other approaches may be required (Huang et al., 2024). For example, Kuhn et al. (2023) use natural language inference to cluster semantically equivalent answers and Tian et al. (2023) use an LLM as a judge of correctness. After consolidating similar events, the relative frequency $f$ of these events is a measure of the LLM’s uncertainty in those events, although this may not be a well-calibrated probability.
3.2 Post-hoc calibration
In general, neural networks are prone to miscalibration, even when trained using proper losses such as the cross-entropy or log-loss. A common remedy is to apply post-hoc calibration methods, which usually involve some form of regression on the predicted scores to transform them into meaningful probabilities. Two common options are isotonic regression and Platt scaling (sometimes called temperature scaling) (Guo et al., 2017). Our approach uses a model’s predictions on $\{X^{\text{cal}},Y^{\text{cal}}\}$ to diagnose and mitigate badly-calibrated initial model probabilities. We fit an isotonic regression model [^5] Implemented using scikit-learn 1.5.2 on our calibration set by comparing the predicted scores to observed labels. [^6] We discuss the effect this has in §missing 6.3. We compare each prediction $\hat{y}$ with score $f$ to observed events $y$. This yields a calibration map $c:\mathbb{R}\rightarrow[0,1]$ we apply to the relative frequencies of events from samples in the previous step to yield probabilities.
3.3 Self-annotation and fine-tuning
We compute the calibrated probability $p=c(f)$ associated with each prediction in the held-out calibration data, and choose a mapping into discrete confidence bins. Several options are possible for this binning function $b$, including adaptive schemes as well as uniform schemes, the number of bins $B$, and so on. In this paper, we focus on a simple fixed-width scheme with $5$ bins. Let $\hat{Y}$ denote the set of all predictions on $X^{\text{cal}}$, and, if the model was previously fine-tuned on a supervised training set $X^{\text{train}}$, we include predictions on $X^{\text{train}}$. We deterministically transform each prediction and calibrated confidence into a training example for a round of supervised fine-tuning by verbalizing the corresponding bin in the answer. For example, the fifth of five bins may correspond to “very high confidence.” The token sequences chosen to encode each bin are arbitrary; for easy interpretability, we use short descriptors in this paper, namely “very low confidence,” “low confidence,” “medium confidence,” “high confidence,” and “very high confidence.”
In our scheme, we simply append the verbalized confidence to all answers. For instance, if the model generates 900 correct answers and 100 incorrect answers, there are two available data points that could potentially be added to the dataset:
<correct answer> (with very high confidence)
<incorrect answer> (with very low confidence)
While correct answers should always be added as training data, appending the confidence scores to incorrect answers may improve the model’s ability to correctly verbalize its own confidence. However, it may also decrease the accuracy of the QA model. We introduce a hyperparameter to control the number of incorrect answers added to the training data. In §missing 6.4, we further discuss the impact of this hyperparameter.
Starting from the sampled model, we perform supervised fine-tuning on these self-annotated targets with verbalized confidences to estimate a second model capable of verbalizing its confidence. If training an instruction-tuned model, we append “Additionally state how confident you are in your answer.” to the preexisting instruction. At inference time, we obtain predictions and verbalized confidences from this new model on held-out test data. We remark that our model incurs little additional cost at inference time, as opposed to other confidence elicitation methods which require inference-time sampling Farquhar et al. (2024); Xiong et al. (2024).
4 Experimental Setup
We examine the efficacy of uncertainty distillation in two settings, with consistent metrics and baseline. In the first, we consider uncertainty quantification on a model that has been fine-tuned to perform a single task. In the second, we consider a multi-task instruction-tuned model on a set of unseen QA tasks, and examine whether our method can achieve reasonable estimates of uncertainty without the model first being fine-tuned on those tasks.
4.1 Supervised fine-tuning experiment
Dataset
We perform these experiments using the SQuAD benchmark (Rajpurkar, 2016). This is a machine-reading task where each question consists of a passage of text and one or more associated questions, each of which is answerable based on the text itself. As the test set has not been publicly released, we use the splits proposed by Du et al. (2017), which divides the publicly available available training and validation splits into train, test, and validation splits. We consider the first 60,000 examples in the training set to be training data, and the remainder to be our calibration set.
Model
We apply uncertainty distillation to T5-base (Raffel et al., 2020) finetuned on a portion of SQUAD. Using a comparatively old and small model ensures that there will be model uncertainty on a reasonably simple dataset like SQUAD. We use defaults for most hyperparameters; in §missing 6, we discuss how uncertainty distillation is impacted by various method-specific hyperparameters.
4.2 Instruction tuning experiment
Dataset
While it is useful to be able to accurately extract confidences from a task-specific language model, in practice modern LLMs are rarely fine-tuned to any specific task. Instead, they are trained to follow instructions such that, in theory, they can perform any task that can be specified in natural language without having necessarily seen that task during training. Thus, it is important that our method can teach a language model to correctly verbalize confidences for a previously unseen task.
To evaluate this, we leverage the Super-NaturalInstructions dataset (SNI; Wang et al., 2022). We select $15$ English Q&A tasks with short-form answers. We focus on Q&A tasks for which a single correct answer exists (e.g. multiple choice problems, short-form span extraction, math problems, etc.) and thus for which correctness of a model’s prediction can reliably and efficiently be computed after normalizing lexical forms without resorting to methods such as LLM verification.
Model
We instruction-tune a T5-Large model on a remaining subset of the English tasks in the SNI dataset, making sure to explicitly hold out the $15$ tasks we use in our uncertainty distillation experiments. The result is an instruction-tuned model which we refer to as Instruct-T5, capable of performing our target Q&A tasks without having seen these tasks during training. In other words, the samples we obtain from this model do not require Instruct-T5 to be pre-trained on that specific task. See Appendix C for more details on our data selection and instruction-tuning. We train and evaluate uncertainty distillation on the combined dataset of these tasks and report the performance over the metrics described in §missing 4.3.
4.3 Metrics
We report the area under the receiver operating characteristic curve (AUROC), [^7] Calculated using scikit-learn 1.5.2 which represents the probability that a randomly chosen correct answer will be in a higher-confidence bin than a randomly chosen incorrect answer. This metric is well established in previous literature (see e.g., Hu et al. (2023)), and compares the relative rather than absolute probabilities, which allows us to use it effectively with discrete verbalized confidences. We plot the percentage of accurate answers in each bin to demonstrate that our model’s predictions correspond well with accuracy. We also report overall model accuracy, to evaluate the tradeoff between accuracy and calibration. Finally, we report accuracy of high-confidence predictions (in “very high” and “high” bins). As an established use-case for verbalized confidences is to reject lower-confidence predictions, this provides information about how useful the LLM’s predictions are. [^8] We do not report Expected Calibration Error (ECE) (Nixon et al., 2019), as it requires comparing a continuous score to the prediction’s true label, while we have five discrete bins without an obvious corresponding numerical score. Furthermore, it requires the choice of several hyperparameters which can have a large impact on performance.
4.4 Baseline
As our tasks are relatively short (with an average of fewer than five tokens in an answer), we report a lexical baseline to compare to our semantic method. [^9] In Appendix A, we also experiment with another baseline based on semantic clustering Kuhn et al. (2023). As it requires sampling 20 answers at inference time, we do not compare it here to methods that generate a single answer. Despite that, we find that it fails to replicate our performance. We extract token-level probabilities from the language model on our training and calibration splits and use this to train an isotonic regression model to calibrate the average token-level probability for each answer. [^10] We use the average probability rather than the sequence probability to normalize over different lengths, as Kuhn et al. (2023) find this improves performance. We find that using this calibrator to map the average calibrated token probability of our test set to a confidence bin (e.g. 90% to “high confidence”) is successful on our tasks, allowing us to compare our semantic method to a strong baseline.
Model Method AUROC Overall Accuracy High Accuracy
T5-base Uncertainty distillation 0.805 0.711 0.852
Lexical baseline 0.771 0.811 0.865
Instruct-T5 Uncertainty distillation 0.751 0.449 0.839
Lexical baseline 0.667 0.387 0.754
Table 1: AUROC and accuracy metrics for T5-base, trained on SQUAD, and Instruct-T5, trained on a combination of factual QA tasks. We find that models trained with uncertainty distillation match or outperform lexical uncertainty in verbalizing confidences on SQUAD-T5 and Instruct-T5.
5 Results and discussion
Figure 2 shows our results comparing uncertainty distillation to the lexical uncertainty baseline in terms of average accuracies in each confidence bin. In plots like this, the goal is to observe a diagonal trend line where outputs reported to have high confidence indeed have high accuracy, and those in the low confidence bins have lower accuracy. In the SQUAD experiment, we find that both uncertainty distillation and the lexical baseline fit the diagonal trend very well, showing that the methods work as intended. In the SNI experiment, the diagonal trend also holds up reasonably well, with minor exceptions at the very high/low bins.
Table 1 summarizes these plots in terms of AUROC score. Our uncertainty distillation achieves AUROC of 0.805 in the T5-base SQUAD experiment, slightly outperforming lexical baseline’s AUROC of 0.771. In the Instruct-T5 SNI experiment, uncertainty distillation achieves 0.771 AUROC compared to 0.667 from lexical baseline. We conclude that uncertainty distillation is an effective alternative to the lexical baseline, even on short-form tasks where lexical uncertainty can be expected to perform well. The advantage of uncertainty distillation in practice is that the language model is able to accurately communicate its own uncertainty without relying on separate models or output token probabilities, matching the lexical baseline without using the same mechanism or inputs. We find that the verbalized confidences produced by uncertainty distillation are highly interpretable, with high correspondence between accuracy of answers within a bin and that bins confidence.
The table also reports the accuracy of the highest confidence bin and the overall accuracy across all bins. While AUROC is the main metric for assessing calibration performance, accuracy is also useful for understanding the nuances of the result. Note that overall accuracy of T5-base on SQUAD dropped from 0.811 to 0.711 after performing uncertainty distillation; on the other hand, overall accuracy of Instruct-T5 on SNI rose from 0.387 to 0.449. Some changes in model accuracy are expected due to the fine-tuning step described in §missing 3.3; we discuss the potential tradeoffs between accuracy and calibration further in §missing 7.
In more detail, for the SNI QA tasks we find that uncertainty distillation results in underconfidence in the lowest bin– the accuracy in the “very low” bin is higher than the ideal range for that bin (between 0 and 20% accuracy). This is likely due to the additional fine-tuning performed in uncertainty distillation. As Instruct-T5 has not been trained on the tasks before sampling, additional fine-tuning improves its performance on these tasks (as can be seen the increased overall accuracy). The model is more likely to get an answer right after fine-tuning, but the annotations do not reflect that change, resulting in underconfidence in those bins. Conversely, the lexical baseline is overconfident in the highest bin, with the accuracy falling well under the ideal range (between 80% and 100% accuracy). This may indicate that token probabilities are a less reliable measure of uncertainty when the language model has not been previously trained on in-distribution training data.
Figure 2: Average accuracy within each confidence bin for our main experiments. We find that our confidence bins are relatively reasonable, with low-confidence bins containing a low proportion of correct answers and high-confidence bins containing a high proportion.
6 Analysis
Model Method AUROC Overall Accuracy High Accuracy
FLAN-T5 Uncertainty distillation 0.873 0.614 0.875
Lexical baseline 0.892 0.657 0.912
Table 2: AUROC and accuracy metrics when using FLAN-T5, which does not have an unseen calibration set. We find that our semantic uncertainty method performs equivalently to a lexical uncertainty method on FLAN-T5.
6.1 Assumption of calibration set
In §missing 3, we assume that we have access to held-out calibration data. However, due to the opacity and size of pretraining datasets, it is increasingly challenging to guarantee that this assumption holds. We test our method on FLAN-T5 Chung et al. (2022), an instruction-tuned model trained on a dataset containing our tasks. Importantly, we not only verify that Flan-T5 has been instruction-tuned on our tasks, but has seen samples from the calibration set of our test tasks. This allows us to investigate the effect of data contamination on calibration of verbalized confidences.
We report our results in Table 2. We find that while uncertainty distillation still produces meaningful confidence bins, it no longer outperforms lexical uncertainty. We conclude that our method works in the absence of this assumption, but not as effectively as token-level probabilities, which are likely well-calibrated due to the model’s previous training on these examples. We discuss results for FLAN-T5 further in §missing 6.3 and §missing 6.4, and find that its behavior differs significantly from results on models where we have an unseen calibration set.
6.2 Number of samples
Our Monte Carlo estimation of probability requires sampling repeatedly from a model before normalizing and calculating probability. In Figure 3, we show that the number of samples used to estimate the initial probabilities has a significant impact if chosen to be too low; however, there are diminishing returns as the number of samples increases. We choose to use 1,000 samples in all of our experiments, as more than that is unlikely to achieve anything but marginal improvement.
Figure 3: Curve showing the AUROC as a function of number of samples on the SQUAD dataset.
6.3 Ablation of post-hoc calibration
If the model’s initial predictions are poorly calibrated, the post-hoc calibration step should help to better align probabilities in the training data with the true likelihood of success. In Table 3, we show the results of models trained with and without this post-hoc calibration step. We find no apparent benefit of post-hoc calibration for T5-base or Instruct-T5. These models are already well-calibrated on their respective domains (see Appendix B), and thus the post-hoc calibration step does not have a significant impact; had the model probabilities been worse-calibrated, the post-hoc calibration step would have increased importance. In the case of FLAN-T5, post-hoc calibration decreases AUROC. Notably, FLAN-T5 does not have unseen calibration data, while Instruct-T5 and T5-base are both trained excluding calibration data. This suggests that in cases when unseen calibration data cannot be obtained, uncertainty distillation may be more effective without the post-hoc calibration step.
Model With Post-hoc No Post-hoc
T5-base 0.804 0.800
Instruct-T5 0.751 0.751
FLAN-T5 0.873 0.883
Table 3: AUROC of models with and without post-hoc calibration at training time. We find that there is no notable performance increase on these datasets with post-hoc calibration.
6.4 Adding incorrect examples
While adding incorrect examples into the training data has the potential to provide more examples at different levels of confidences, it also is likely to increase the likelihood that a model generates an incorrect answer. To demonstrate this effect, in Table 4, we show the AUROC and accuracy for models trained with different amounts of incorrect samples. For T5-base and Instruct-T5, we find that adding only two incorrect samples per correct sample dramatically increases AUROC while decreasing accuracy. While this would seem to indicate a fundamental tradeoff between accuracy and calibration, we find that the same is not as obviously true for FLAN-T5; while the accuracy may decrease and AUROC may increase, the effects are not as significant as they are for T5-base or Instruct-T5. One possible interpretation of this is that its predictions are shaped by the fact that the data was included in its instruction-tuning corpus, leading to less dramatic shifts when trained.
While adding incorrect samples may improve AUROC, it the number of training examples by a factor of the number of incorrect examples added (e.g. a training set with 100 examples would train on 100 augmented answers with 0 incorrect examples added, 200 augmented answers with one incorrect example added, etc.) This leads to increased compute at training time. For this reason, in addition to the decreased accuracy, we recommend adding a low number of incorrect examples to the training dataset, and in our experiments limit to two incorrect answers per question.
0 1 2 3
T5-Base
AUROC 0.707 - 0.805 -
Accuracy 0.791 - 0.711 -
Instruct-T5
AUROC 0.723 - 0.751 -
Accuracy 0.529 - 0.449 -
FLAN-T5
AUROC 0.868 0.876 0.873 0.883
Accuracy 0.609 0.620 0.614 0.611
Table 4: AUROC of models trained with varying numbers of incorrect examples allowed per question. There is a general trend towards increasing AUROC and decreasing accuracy when incorrect examples are included, although this is less pronounced for FLAN-T5. FLAN-T5 displays only a slight version of this trend; we evaluate over more points to demonstrate that, while slight, the same phenomenon occurs.
7 Conclusion
Findings
We find considerable evidence that using uncertainty distillation to verbalize semantic uncertainty can perform comparably to or better than lexical uncertainty in QA tasks, despite operating using an entirely different mechanism. We demonstrate that our method is successful in all three settings we test it in– a small fine-tuned model, a large instruction-tuned model with unseen data, and an instruction-tuned model calibrated with information in its instruction-tuning set. We also find that in cases where the initial model has not been tuned on the specific task before uncertainty quantification methods are used, the lexical uncertainty baseline tends towards overconfidence, while uncertainty distillation tends to be underconfident.
One high-level takeaway is that there appears to be a tradeoff between an LLM’s ability to predict its own confidence and overall model accuracy. In our experiments, interventions that improve AUROC decrease accuracy and vice versa. While some amount of this may be remedied by increased hyperparameter tuning or small operations, it is also plausible that there are some fundamental tradeoffs.
Future work
In this work, we demonstrate that uncertainty distillation exhibits strong performance comparable to a lexical baseline only on tasks where the lexical method succeeds. Future work should examine its effectiveness in cases where a lexical baseline fails. We focus on strings that are quickly and easily normalized with preprocessing functions, but our method could easily be applied to tasks outside simple QA through the use of LLM verifiers to calculate binary correctness. Future work may also investigate the robustness of the model’s internal representation of uncertainty to domain shifts, particularly in the case of multi-task generalization—do the confidence scores remain interpretable and well-calibrated on similar unseen tasks without further training?
Limitations
While we anticipate that increasing model size would only increase our method’s efficacy, we do not test on models larger than T5-large, both for computational reasons and because we cannot ascertain which data is in training sets for many larger models. We also run on relatively simple QA tasks, where computing uncertainty is comparatively straightforward. Our method is also not infallible, with imperfect accuracies in high-confidence bins– incorrect high confidence predictions may present a small risk of prompting users to trust predictions more than they otherwise would have.
Appendix A Comparison to semantic clustering
In Table 5, we compare uncertainty distillation to semantic clustering as described by Kuhn et al. (2023). In this method, twenty answers are sampled at inference time, and DeBERTaHe et al. (2020) is used to determine how many are semantically equivalent; the most common answer is evaluated, and the proportion of answers in that cluster is considered to be the probability. Similarly to our own method and the baseline, we evaluate AUROC on the binned scores. We find that semantic clustering fails to outperform uncertainty distillation despite using 20 times the compute, and additionally find that overall accuracy with semantic clustering is highly variable, though accuracy in the top bins remains high.
Model Method AUROC Overall Accuracy High Accuracy
Instruct-T5 Uncertainty distillation 0.751 0.449 0.839
Semantic clustering 0.708 0.65 0.917
FLAN-T5 Uncertainty distillation 0.873 0.614 0.875
Semantic clustering 0.868 0.346 0.879
Table 5: AUROC and accuracy metrics compared to semantic clustering. We find that uncertainty distillation compares favorably with a single sample, rather than the 20 samples required at inference time by semantic clustering.
Appendix B Model calibration on QA tasks
While LLM overconfidence is well-documented on many tasks, we find that our LLMs are well-calibrated on the short QA tasks we use for uncertainty estimation. Figure 4 shows the reliability diagrams for T5-base on SQUAD and Instruct-T5 on SNI. The models’ predicted confidences align well with their actual accuracies; a post-hoc calibrator will not significantly alter the output probabilities.
Figure 4: Initial calibration of our T5-base and Instruct-T5 model. Both models are well-calibrated in their respective domains, indicating that post-hoc calibration may not be necessary.
Appendix C SuperNatural-Instructions Tasks
C.1 Target Calibration Tasks
As we describe in §missing 3, in this work we rely on the assumption that our target-tasks have a correct answer, in the sense that it can be easily verified that an answer is right or wrong. Although this is not a strict necessity for calibration, it allows for us to define our buckets in terms of expected accuracy, rather than e.g. an expected score. We therefore focus on short-form Q&A tasks, question-answer pairs whose answers consist of either selection from a fixed answer set (e.g. multiple choice or fixed choice) or single-word answers. We identify $15$ tasks from the SuperNatural-Instructions dataset (Wang et al., 2022) that fit our criteria, and hold out these tasks as our uncertainty prediction tasks.
These tasks are split across $4$ rough task types: Multiple Choice tasks involve selecting an answer from a set of choices, where the response is either a number or letter indicating the choice; Fixed Choice tasks involve selecting an answer from a pre-defined set of choices that are constant across the task (e.g. respond with either True or False); Span Selection tasks involve selecting the correct span of text from context and responding with that span as the answer; Open Answer involves generating the answer to the question in an open-ended way, i.e. the answer is not provided in the context.
For all tasks, we ensure that the answers are no more than 2 words long, making it easy to perform normalization and verify accuracy for each question. The tasks are shown in Table 6; for each task, we use $10\%$ of the samples as a validation set, $10\%$ of the samples as a held-out test set, use the remaining $80\%$ of the data to form our calibration set.
Task Type Task Name
Multiple Choice task580-socialiqa-answer-generation
task309-race-answer-generation
task1297-qasc-question-answering
task1420-mathqa-general
task228-arc-answer-generation-easy
task1286-openbookqa-question-answering
task1431-head-qa-answer-generation
task1731-quartz-question-answering
task750-aqua-multiple-choice-answering
Fixed Choice task380-boolq-yes-no-question
task1661-super-glue-classification
Span Selection task002-quoref-answer-generation
task041-qasc-answer-generation
Open Answer task591-sciq-answer-generation
task898-freebase-qa-answer-generation
Table 6: The tasks and task types that we select from the SuperNatural-Instructions dataset for validating and testing our calibration method.
C.2 Instruction-Tuning Tasks
Because most modern instruction-tuned models are trained on all of Super-NaturalInstructions, they have seen the our calibration target tasks during instruction-tuning. Therefore, we instruction-tune our own T5 model to test the effectiveness of our method on unseen tasks. Our model is trained on a subset of the SuperNatural-Instructions dataset (Wang et al., 2022). Specifically, we instruction-tune on the English split used in the original paper but we take out our target calibration tasks identified in §missing C.1. This gives us a training dataset of $879$ instruction-tuning tasks, with a total of roughly 1.2M training samples total.
To validate our models instruction-following capabilities, we use the in-context learning test set from SuperNatural-Instructions, which contains 95 additional held out tasks from task categories that are not seen in the training dataset.
Appendix D Instruction-Tuning T5
We follow a standard recipe for instruction-tuning T5-Large, established in Wang et al. (2022). Specifically, we tune the model for 3 epochs with a batch size of $16$ and a learning rate of $5\times 10^{-3}$. We use the AdamW optimizer, and a constant learning rate schedule after a warmup period of $500$ steps. During instruction-tuning, we train the model with the semantic definition of each task prepended to the task input, and we similarly prompt the model when performing our target Q&A tasks.
Appendix E Resource Reporting
E.1 Compute Resources
Here we report the compute resources used in this work. Instruction-tuning T5 took a total of $200$ GPU hours across $4$ NVIDIA-V100s. Running uncertainty distillation on Instruct-T5 and FLAN-T5 took 16 hours per model on a single NVIDIA-H100. Finetuning T5-base on SQUAD for our initial model took 3 hours on a single NVIDIA RTX 2080, and training using uncertainty distillation took 8 hours on a single NVIDIA-V100. Our lexical baseline for SQUAD took one hour on one NVIDIA RTX 2080, and for SNI took three hours on one NVIDIA RTX 2080. Sampling for SQUAD took a total of 60 GPU hours on NVIDIA-V100s, and for SNI took 45 GPU hours on NVIDIA-A100s.
E.2 Resource Intended Use
Super-NaturalInstructions (SNI) is an open-source instruction tuning dataset, released under the Apache License. [^11] Available here: https://github.com/allenai/natural-instructions The intended use of SNI is to instruction-tune language models to learn to follow instructions, and to evaluate a model’s ability to follow instructions on unseen tasks. While we use the SNI dataset for precisely this purpose during instruction-tuning, we also use $15$ held-out tasks to serve as uncertainty quantification tasks. This does not necessarily fall under the intended use of instruction-tuning; however, the authors of SNI also mention that the dataset may serve as a large, multi-task natural language resource (Wang et al., 2022), and our usage of the target calibration tasks does fall under this use case.
The Stanford Question Answering Dataset (SQUAD) Rajpurkar (2016) is distributed under the Creative Commons Attribution-Sharealike 4.0 license, which permits use of the dataset as long as it is properly attributed and as long as the results are distributed under the same license. As we cite the paper and plan to publically release our code and models after acceptance, our use of this dataset is permitted under this license.
Appendix F Uncertainty distillation hyperparameters
In Table 7, we show the training hyperparameters for uncertainty distillation training.
Model Epochs Learning rate Batch size Grad accumulation steps
T5-base (initial) 1 3e-5 12 1
T5-base (Uncertainty distillation) 3 3e-5 12 1
Instruct-T5 (Uncertainty distillation) 3 3e-5 1 32
FLAN-T5 (Uncertainty distillation) 3 3e-5 1 32
Table 7: Hyperparameters for training all models but Instruct-T5 (see Appendix D for details). All models are trained with the AdamW optimizer.
Appendix G Algorithm
Algorithm 1 Uncertainty distillation
Language model $f_{\theta}$ with params $\theta_{0}$
Calibration set $S^{cal}=\{X^{cal},Y^{cal}\}$
$S^{scored}\leftarrow\emptyset$
for $(x,y)\in S^{cal}$ do
     $D\leftarrow\{\hat{y}_{i}\}_{i=1}^{N}\sim f_{\theta}(x)$
     Normalize $D$ by semantics, and count
     for $\hat{y}\in D$ with count $n$ do
         $f\leftarrow\frac{n}{N}$
         $S^{scored}\leftarrow S^{scored}\cup\{(x,\hat{y},y,f)\}$
     end for
end for
$c()\leftarrow$ isotonic_regression$(S^{scored})$
$S^{vc}=\emptyset$
for $(x,\hat{y},y,f)\in S^{scored}$ do
     if filter$(\hat{y},y)$ then
         continue
     end if
     $p\leftarrow c(f)$
     $b\leftarrow$ bin$(p)$
     $z\leftarrow$ verbalize_confidence_map$(\hat{y},b)$
     $S^{vc}\leftarrow S^{vc}\cup\{(x,z)\}$
end for
$\mathcal{L}(\theta)\leftarrow\mathbb{E}_{(x,z)\in S^{vc}}[NLL(f_{\theta}(x),z)]$
$\theta_{cal}\leftarrow$ train$(\theta_{0},\mathcal{L})$
Return $\theta_{cal}$